OpenStudy (anonymous):

I am not sure if i understand how to add fractions with unlike denominators-is this correct: 1/6+5/9= 1/9 ? 0r 3/27

OpenStudy (anonymous):

1/6+5/9= (1/6)(3/3) + (5/9)(2/2) = 3/18 + 10/18 = 13/18

OpenStudy (anonymous):

neither. Change to common denominator \[1/6*3/3+5/9*2/2=3/18+10/18=13/18\]